OpenStudy (anonymous):

You would like to have $170,000 in 12 years by making regular deposits at the end of each month in an annuity that pays 8.25% compounded monthly. How much of the $170,000 comes from interest? Round your answer to the nearest dollar.

OpenStudy (tkhunny):

What's your plan? I would build it from basic principles. i = 0.0825 -- Annual Interest Rate j = I/12 = 0.006875 -- Monthly Interest Rate r = 1+j = 1.006875 -- Monthly Accumulation Factor Last Payment: Pmt*1 2nd to last payment: Pmt*r 3rd to last payment: Pmt*r^2 etc All told: \(Pmt\cdot(1 + r + r^{2} + ... + r^{143}) = 170000\) Can you add those up inside the parentheses?